OpenStudy (anonymous):

At the Boston Children's Museum, it was estimated 38% of all visitors are from in state. On Saturday, 8,000 people attended the museum. What is the number of out of state people in attendance?

OpenStudy (anonymous):

4960

OpenStudy (anonymous):

\[8000(\frac{100-38}{100})=4960\]

OpenStudy (anonymous):

your equation looks nice i can never get the division thing to work like that 8(

OpenStudy (anonymous):

\[8000(100-38/100)=4960\]